OpenStudy (anonymous):

Paula ran for 195 minutes this weekend. She ran 25 minutes more on Saturday than Sunday. How long did she run each day?

OpenStudy (anonymous):

set x = number of minutes she ran on saturday, then sunday would be x - 25 since "She ran 25 minutes more on Saturday than Sunday". and u know she ran 195 minutes in total, saturday plus sunday which is x + (x - 25), which simplifies down to 2x-25 = 195, add 25 from both sides u get 2x = 220, and u know x = 110, which is the number of minutes she ran on saturday, and we now sunday = x-25, which is 85. so she ran 110 minutes on Saturday and 85 minutes on Sunday. 110 + 85 = 195 so it is correct